# Baltimore Atlas: FreqWeaver Adapter for Semi-supervised Ultra-high Spatial Resolution Land Cover Classification

### v1

@@ Abstract.
Ultra-high Spatial Resolution Land Cover Classification is essential for fine-grained land cover analysis, yet it remains challenging due to the high cost of pixel-level annotations, significant scale variation, and the limited adaptability of large-scale vision models. Existing methods typically focus on 1-meter spatial resolution imagery and rely heavily on annotated data, whereas practical applications often require processing higher-resolution imagery under weak supervision. To address this, we propose a parameter-efficient semi-supervised segmentation framework for 0.3 m spatial resolution imagery, which leverages the knowledge of SAM2 and introduces a remote sensing–specific FreqWeaver Adapter to enhance fine-grained detail modeling while maintaining a lightweight design at only 5.96% of the total model parameters. By effectively leveraging unlabeled data and maintaining minimal parameter overhead, the proposed method delivers robust segmentation results with superior structural consistency, achieving a 1.78% improvement over existing parameter-efficient tuning strategies and a 3.44% gain compared to state-of-the-art high-resolution remote sensing segmentation approaches.
High-resolution remote sensing, Semi-supervised segmentation, Large model priors, Small object segmentation, Structural consistency
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Computing methodologies Image segmentationccs: Computing methodologies Semi-supervised learningccs: Computing methodologies Transfer learningccs: Computing methodologies Convolutional neural networks
@@ 1. Introduction
High-precision segmentation of remote sensing (RS) imagery serves as a cornerstone in a wide range of real-world applications, including land cover classification, urban planning, environmental monitoring, and disaster response (Lv et al., 2023; Guo et al., 2024). With the advent of ultra-high-resolution (UHR) imagery—particularly at the 0.3-meter ground sampling distance—RS data now enables significantly richer spatial detail, allowing for the identification of fine-scale urban features such as vehicles, street furniture, residential swimming pools, and narrow alleyways (Gui et al., 2024). This spatial granularity facilitates parcel-level land use analysis and fine-grained urban modeling, thereby unlocking new frontiers for RS research and practical applications (Li et al., 2024).
However, the transition to UHR imagery introduces substantial challenges for semantic segmentation models. While existing approaches demonstrate robust performance on datasets with 1-meter resolution, their effectiveness often deteriorates when applied to finer resolutions such as 0.3 meters (Mendieta et al., 2024; Ma et al., 2025; Zhang et al., 2022). The increased level of detail exacerbates the complexity of capturing fine-grained structural boundaries and contextually diverse object classes (Huang et al., 2025). Effective segmentation at this scale requires models capable of simultaneously modeling local geometry and global semantics, motivating the development of novel architectures tailored to the intricacies of UHR data.
Moreover, a major bottleneck in the advancement of ultra-high-resolution (UHR) RS segmentation lies in the reliance on large-scale, pixel-wise annotated datasets for fully supervised training (Liu and Zhang, 2025). While effective under ideal settings, manual annotation of ultra-high-resolution RS imagery is prohibitively time-consuming and labor-intensive, limiting the scalability and practical deployment of such models (He et al., 2024). Moreover, when operating under limited supervision, these models often exhibit performance degradation and poor generalization across diverse scenes and geographies (Zhang et al., 2024).
Recently, the emergence of pre-trained foundation models such as the Segment Anything Model (SAM) (Kirillov et al., 2023a; Ravi and Gabeur, 2024) has demonstrated strong generalization capabilities across diverse visual tasks. Trained on extensive image-mask pairs, SAM2 (Ravi and Gabeur, 2024) encapsulate transferable semantic priors and offer a promising path toward alleviating annotation dependency via transfer learning on downstream tasks like high-resolution RS image segmentation (Zhang et al., 2025).
Although full fine-tuning of foundation models can enhance performance, it also incurs substantial computational and memory overheads (Han et al., 2024). Parameter-Efficient Fine-Tuning (PEFT) methods, such as Adapters (Houlsby and Giurgiu, 2019) and Low-Rank Adaptation (LoRA) (Hu and Shen, 2022), have emerged as viable alternatives for resource-conscious adaptation. However, due to the pronounced domain gap between the natural or medical images used for pre-training and the unique characteristics of RS data, conventional PEFT strategies often fail to generalize well in RS scenarios (Khanna and Irgau, 2024; Stojnic and Klemmer, 2025). Moreover, remote sensing imagery encompasses a wide spectrum of object scales—from large-scale vegetated regions to fine-grained human-made structures—necessitating robust multi-scale feature modeling capabilities (Cheng and Lei, 2023).
To this end, we propose FreqWeaver, a frequency-temporal modulation adapter designed to enhance the adaptability of foundation models like SAM2 to ultra-high-resolution (UHR) remote sensing imagery. FreqWeaver decomposes input features into low- and high-frequency components via spectral filtering, which are then processed using large and small receptive field convolutions to respectively capture global semantics and fine-grained structures. A spatial-domain residual branch is further incorporated to complement missing local details. This design not only leverages the semantic priors encoded in SAM2 for basic object recognition, but also enhances the model’s capacity to represent diverse spatial patterns in UHR data. In particular, the proposed multi-scale feature interaction mechanism significantly improves the model’s ability to delineate fine boundary details of large-scale objects and accurately identify small, morphologically complex structures—thereby enabling precise and robust segmentation across heterogeneous geographic scenes under strict parameter constraints.
Remote sensing (RS) systems, owing to their extensive spatial coverage and high temporal resolution, continuously collect vast amounts of unlabeled imagery, offering a rich yet underutilized resource for semi-supervised learning (SSL) (Yan and Yang, 2024). To address the challenges posed by limited supervision and high annotation costs in high-resolution RS segmentation, we propose a semi-supervised fine-tuning framework built upon a pre-trained foundation model. The framework integrates a teacher–student consistency learning paradigm with an uncertainty-aware weighting strategy, enabling efficient and robust domain adaptation. By dynamically modulating the contribution of pseudo-labels during training, the method effectively suppresses the adverse impact of noisy supervision—an issue particularly prevalent in RS imagery due to complex backgrounds, ambiguous boundaries, and high inter-class similarity. As a result, it significantly enhances model robustness and generalization across diverse geographic regions, acquisition conditions, and sensor types (Fan and Zhou, 2025), while simultaneously reducing reliance on costly manual annotations.
@@
  • PEFT-Driven Semi-Supervised Tuning of SAM2: We propose an Adapter-based PEFT strategy integrated with an uncertainty-guided consistency regularization mechanism for the SAM2 model (Ravi and Gabeur, 2024). This approach significantly reduces the dependency on labeled data, achieving efficient segmentation performance while updating fewer than $5.96\%$ of the total model parameters.
  • 0.3m-Resolution Remote Sensing Dataset for Fine-Grained Segmentation: To support detailed land cover analysis, we construct a high-resolution (0.3m) remote sensing dataset over Baltimore. Compared to conventional 1m-Resolution datasets, it offers finer spatial granularity and greater object diversity, providing a more challenging and representative benchmark for evaluating segmentation models under real-world urban conditions.
  • FreqWeaver Adapter for Disentangled Frequency–Temporal Multi-Scale Modeling: To better capture the fine-grained structures and scale-diverse patterns in high-resolution RS imagery, we introduce FreqWeaver, a frequency–temporal modulation adapter that decomposes features into low- and high-frequency components for separate processing via large and small receptive field convolutions. A complementary spatial-domain branch further enriches semantic representation. This design enhances SAM2’s capacity to model both global context and local detail, enabling more precise and parameter-efficient multi-scale segmentation.
  • Uncertainty-Aware Semi-Supervised Learning: Considering the high error rates of pseudo-labels in remote sensing imagery—caused by complex backgrounds, fine-grained boundaries, and inter-class spectral similarity—we propose an uncertainty-aware semi-supervised framework that downweights unreliable predictions based on entropy. This design effectively suppresses misleading gradients from noisy supervision and improves the robustness of segmentation across diverse regions, sensors, and seasonal conditions.
@@ 2. Relate Work
2.1. Foundation Models in Remote Sensing
Foundation Models (FMs), notably Vision Transformer (ViT)-based architectures (Dosovitskiy et al., 2021), learn transferable representations from extensive unlabeled data, enabling strong generalization across diverse tasks. Adapting FMs to remote sensing (RS) is challenging due to RS imagery’s distinct characteristics (multi-scale spatial structures, rich spectral content, complex semantic distributions) that diverge significantly from natural images.
To address these challenges, Geospatial Foundation Models (GeoFMs) have emerged. Models like Clay (Model, 2024) and Prithvi (Szwarcman et al., 2025), pre-trained on extensive satellite/aerial data, are optimized for Earth Observation (EO). Recent advances such as S2MAE (spatial-spectral pretraining) (Placeholder et al., 2024b), GeoChat (vision-language grounding) (Placeholder et al., 2024a), and SkySense (multi-modal integration) (Placeholder et al., 2024c) highlight the trend towards task-adaptive geospatial architectures.
General-purpose FMs like the Segment Anything Model (SAM) (Kirillov et al., 2023b) and SAM2 (Ravi and Gabeur, 2024) offer prompt-driven, class-agnostic segmentation. However, the original SAM performs suboptimally on RS imagery, showing limited semantic awareness, difficulty with small/dense objects, and inconsistency in tasks like semantic change detection. These issues stem from its natural image bias and lack of RS domain priors.
To address these SAM limitations, STSAMNet (Yang et al., 2025) integrates unsupervised domain adaptation (UDA), pseudo-label refinement, and adapter layers for improved building extraction. SCNet (Jiang et al., 2025) uses multi-scale feature fusion and semantic alignment to adapt SAM2 for semantic change detection. Prompt generation and CLIP-based supervision are also explored for semantic guidance in RS.
These adaptations show a shift from universal zero-shot inference to domain-specialized fine-tuning. Thus, RS is not just adopting FMs but also shaping their development. This is driven by RS challenges like scale variation, semantic ambiguity, and the need for label-efficient geospatial analysis.
2.2. Semantic Segmentation for Complex Remote Sensing
Remote sensing image segmentation faces unique challenges. Pixel-level annotation scarcity is primary, severely limiting fully supervised methods. Second, inherent multi-scale spatial heterogeneity necessitates models adept at handling drastic scale variations (small objects to vast regions). Furthermore, high inter-class similarity and significant intra-class diversity (from varying illumination, seasons, and sensors) severely test model feature discriminability.
Semi-Supervised Semantic Segmentation (SSSS) addresses annotation scarcity by leveraging abundant unlabeled data. Key SSSS strategies include consistency regularization and pseudo-labeling(Tarvainen and Valpola, 2018; Chen et al., 2021; Sohn and Berthelot, 2020). For instance, MUCA (Wang et al., 2025) uses a Multi-scale Uncertainty Consistency (MSUC) module and cross-teacher-student attention to address multi-scale objects and inter-class similarity, enhancing multi-layer semantic consistency and discriminability. SegKC (Than et al., 2025) improves pseudo-label quality and efficiency via heterogeneous dual-network co-training with bi-directional knowledge transfer, using a lightweight network for efficient inference.
Unsupervised Domain Adaptation (UDA) addresses domain shifts in remote sensing data, often from varying sensors or locations. For example, STSAMNet (Yang et al., 2025) combines UDA with the Segment Anything Model (SAM) via an adapter, optimizing segmentation for target-domain high-resolution imagery and mitigating performance drops in new regions.
While SSSS and UDA methods partially address label scarcity and domain shifts, accurately segmenting complex spatial structures and multi-scale targets remains a bottleneck. Many methods struggle to capture diverse multi-scale features or introduce artifacts during pseudo-labeling and domain adaptation, especially at boundaries or in textured areas. To overcome these limitations, we explore parameter-efficient solutions that not only reduce reliance on labeled data but also enhance multi-scale representation under minimal supervision, as detailed next.
2.3. Parameter-Efficient Fine-Tuning Method
As foundation models (FMs), including geospatial variants (GeoFMs), continue to grow in size and complexity, full fine-tuning has become increasingly impractical due to high computational costs and risks such as overfitting and catastrophic forgetting. Parameter-Efficient Fine-Tuning (PEFT) offers a scalable alternative by adapting models through the selective update of a small subset of parameters, while keeping the majority of pre-trained weights frozen.
Several PEFT techniques have been proposed to balance efficiency and performance. Adapter tuning, which inserts lightweight modules into transformer layers, has been successfully applied in models like STSAMNet (Yang et al., 2025) for adapting SAM to high-resolution remote sensing. Low-Rank Adaptation (LoRA) (Hu et al., 2021) reduces parameter overhead by constraining updates to low-rank subspaces. Prompt tuning (Jia et al., 2022) and prefix tuning (Li and Liang, 2021) modify task-specific behavior through learned embeddings, without altering the base model. These methods significantly reduce training cost and memory usage, and often enhance generalization by preserving pre-trained representations.
PEFT has gained traction in remote sensing. (Marti-Escofet et al., 2025) benchmarked multiple PEFT methods across diverse GeoFM architectures and Earth observation datasets, showing that PEFT can match or surpass full fine-tuning with far lower resource requirements. (He et al., 2023) introduced Sensitivity-aware PEFT (SPT), which identifies and prioritizes task-relevant parameters for adaptation, integrating LoRA or adapters accordingly. Domain-specific approaches such as AiRs (Hu et al., 2024b), TEA (Hu et al., 2024a), and Earth-Adapter (Hu et al., 2025) further enhance transformer adaptation for Earth observation via residual-enhanced adapters.
@@ 2.3. Parameter-Efficient Fine-Tuning Method
To address key challenges in remote sensing imagery—such as large parameter overhead, aliasing artifacts, and significant object-scale variability—we propose a novel adapter architecture that combines frequency-domain decomposition with specialized convolutional processing. By disentangling high- and low-frequency components through dedicated convolutional branches and incorporating temporal information, the proposed method enables effective multi-scale adaptation while maintaining minimal parameter cost.
Figure 1. Overview of the proposed framework.

### v2

@@ Abstract.
Ultra-high Spatial Resolution (UHSR) Land Cover Classification is increasingly important for urban analysis, enabling fine-scale planning, ecological monitoring, and infrastructure management. It identifies land cover types on sub-meter remote sensing imagery, capturing details such as building outlines, road networks, and distinct boundaries. However, most existing methods focus on 1 m imagery and rely heavily on large-scale annotations, while UHSR data remain scarce and difficult to annotate, limiting practical applicability. To address these challenges, we introduce Baltimore Atlas, a UHSR land cover classification framework that reduces reliance on large-scale training data and delivers high-accuracy results. Baltimore Atlas builds on three key ideas: (1) Baltimore Atlas Dataset, a 0.3 m resolution dataset based on aerial imagery of Baltimore City; (2) FreqWeaver Adapter, a parameter-efficient adapter that transfers SAM2 to this domain, leveraging foundation model knowledge to reduce training data needs while enabling fine-grained detail and structural modeling; (3) Uncertainty-Aware Teacher Student Framework, a semi-supervised framework that exploits unlabeled data to further reduce training dependence and improve generalization across diverse scenes. Using only 5.96% of total model parameters, our approach achieves a 1.78% IoU improvement over existing parameter-efficient tuning strategies and a 3.44% IoU gain compared to state-of-the-art high-resolution remote sensing segmentation methods on the Baltimore Atlas Dataset.
Ultra-high Spatial Resolution Land Cover Classification, Parameter-Efficient Adapter, Semi-Supervised Segmentation, Transfer Learning
journalyear: 2025copyright: ccconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762724isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Computer visionccs: Computing methodologies Image segmentationccs: Computing methodologies Semi-supervised learning settingsccs: Information systems Geographic information systemsccs: Theory of computation Design and analysis of algorithms
@@ 1. Introduction
Ultra-high Spatial Resolution (UHSR) remote sensing imagery, typically at sub-meter scales, offers rich spatial detail that allows accurate identification of individual ground objects and preservation of their structural features (Lv et al., 2023). Compared with coarser imagery, it provides sharper boundaries, finer textures, and the ability to capture small, heterogeneous features that are essential for detailed geospatial analysis (Guo et al., 2024). Leveraging these strengths, UHSR land cover classification converts high-detail imagery into thematic maps that combine broad spatial patterns with precise object-level information (Gui et al., 2024). The fine granularity of UHSR data reduces mixed-pixel effects, improves boundary accuracy, and enables reliable detection of subtle features such as narrow roads, isolated vegetation patches, and variations in roof materials or tree canopy structures. These advantages produce classification results that more closely match real-world surface conditions, supporting parcel-level land use mapping, high-precision urban modeling, ecosystem monitoring, and disaster impact assessment (Li et al., 2024a). By linking large-scale landscape observation with fine-scale object analysis, UHSR-based land cover classification has become an essential tool for modern geospatial applications.
Despite its importance, most current remote sensing segmentation methods still focus on imagery with 1-meter spatial resolution, which is sufficient for identifying coarse land cover categories such as large vegetation patches, water bodies, and major transportation corridors, but struggles to capture fine-grained details such as narrow alleyways, small rooftop structures, or individual vehicles (Mendieta et al., 2024; Ma et al., 2025; Zhang et al., 2022). UHSR imagery, by contrast, enables the recognition of such detailed structures and subtle boundaries, offering the potential for parcel-level urban analysis and more precise land cover mapping (Huang et al., 2025). However, annotated UHSR datasets remain scarce and are extremely time- and labor-intensive to produce (Liu and Zhang, 2025; He et al., 2024), as pixel-level labeling at this resolution demands careful delineation of numerous small and heterogeneous objects. This annotation bottleneck has limited large-scale exploration of UHSR land cover classification in both research and practical applications.
To address these limitations, we propose Baltimore Atlas, a novel framework for UHSR land cover classification. Baltimore Atlas performs multi-scale frequency-decoupled modeling for complex UHSR data patterns, preserving global boundaries while capturing fine-grained local details. Moreover, the framework leverages the prior knowledge of foundation models and integrates a semi-supervised learning strategy to reduce reliance on large annotated datasets. Consequently, the framework achieves high-accuracy and robust UHSR land cover mapping under limited training data availability.
At the heart of Baltimore Atlas is the FreqWeaver Adapter, a frequency-modulated module designed to adapt SAM2 for UHSR imagery. To address the diverse spatial patterns in UHSR data, the module first applies spectral filtering to separate input features into low- and high-frequency components. The low-frequency branch employs large-receptive-field convolutions to capture global semantics, while the high-frequency branch applies small-receptive-field convolutions to preserve fine-grained structures. In parallel, a complementary spatial-domain residual branch restores local details. This design introduces only two additional convolutional kernels, maintaining a parameter-efficient architecture, and enables the model to effectively integrate global context with fine-grained detail for more accurate UHSR land cover classification.
Further, we introduce the Uncertainty-Aware Teacher Student Framework, a semi-supervised framework for UHSR data. The teacher model takes weakly perturbed inputs and generates predictions. We compute pixel-wise entropy from these predictions, and lower entropy means higher confidence. For each pixel, the class with the highest confidence is chosen as the pseudo-label. High-confidence regions give stronger supervision to the student model, which is trained on strongly perturbed inputs. Focusing on these reliable regions helps the framework reduce noise from low-confidence areas, learn more stable representations, and improve performance and generalization in complex scenes. In addition, leveraging large-scale unlabeled data further reduces the reliance on annotated training samples.
@@
  • We propose Baltimore Atlas, a novel framework for UHSR land cover classification that achieves high-accuracy and robust mapping even with limited training data.
  • We design a unified architecture with two core components: FreqWeaver Adapter, which parameter-efficiently adapts SAM2 to complex UHSR data patterns, and Uncertainty-Aware Teacher Student Framework, which uses entropy-based confidence to learn stable representations and reduce complex noise.
  • We introduce the Baltimore Atlas Dataset, a 0.3 m resolution dataset based on aerial imagery of Baltimore City. Experiments on this dataset demonstrate consistent improvements over prior methods.
@@ 2. Relate Work
2.1. Ultra-high Spatial Resolution Land Cover Classification
Early land cover classification mainly used pixel-based methods, treating each pixel as an independent unit classified only by its spectral features. In high spatial resolution imagery, these methods often cause confusion and fragmentation due to the high spectral similarity among different land covers (Roy et al., 2018; Han et al., 2020), leading to the development of Geographic Object-Based Image Analysis (GEOBIA/OBIA) (Chen et al., 2018; Hay and Castilla, 2008; Kucharczyk et al., 2020). GEOBIA aggregates similar pixels into image objects and combines segmentation, feature extraction, and semantic rules to improve classification, but large-scale mapping still requires considerable expert knowledge and manual intervention. To reduce manual effort and improve automation, deep learning has been widely applied to fine-grained land cover classification. LSHR-LCM (Robinson et al., 2019) evaluated multiple deep learning models on 1 m resolution data across the United States, demonstrating the feasibility of convolutional networks for nationwide mapping but showing strong dependence on high-quality labeled data. ABCNet (Li et al., 2021) improved network design to preserve fine spatial details while capturing global context, enhancing both accuracy and stability in high-resolution semantic segmentation. Urban Watch (Zhang et al., 2022) refined the impervious surface class, resulting in a total of nine urban categories, to address material diversity and overly coarse definitions, and incorporated semi-supervised learning to reduce annotation requirements. LEG (Li et al., 2024b) used 30 m low-resolution labels for weak supervision, improving 1 m mapping accuracy and demonstrating the potential for cross-resolution transfer. TLULC (Khan et al., 2024) applied large-scale Transformer pretraining and transfer learning, further enhancing the generalization and robustness of high-resolution mapping. Despite these advances, most methods focus on 1 m or coarser resolutions, leaving UHSR land cover classification largely unexplored.
To address this gap, we propose the Baltimore Atlas, a framework for UHSR(0.3 m resolution) land cover classification that combines tailored transfer learning with semi-supervised learning to adapt to the complex spatial patterns of UHSR imagery, achieving improved performance while reducing reliance on large-scale annotated datasets.
2.2. Parameter-Efficient Adaptation for Multi Scale Remote Sensing Understanding
Foundation models (FMs) such as SAM2 (Ravi et al., 2024) are trained on large-scale natural image datasets with segmentation labels, enabling strong generalization across segmentation tasks. DED-SAM (Qiu et al., 2025) demonstrates SAM2’s zero-shot segmentation capability on remote sensing (RS) imagery. To better align model representations with RS characteristics, GeoFMs such as Clay (Model, 2024), Prithvi (Szwarcman et al., 2024), and S2MAE (Li et al., 2024c) incorporate RS-specific priors through pretraining on satellite and aerial imagery. However, these models still lack task-specific priors for segmentation. As a result, both general-purpose FMs and GeoFMs require adaptation to downstream datasets to achieve strong performance in tasks like land cover classification. Parameter-efficient fine-tuning (PEFT) addresses this need by reducing trainable parameters while maintaining accuracy: LoRA (Hu et al., 2022) constrains weight updates to low-rank subspaces, prompt tuning (Jia et al., 2022) and prefix tuning (Li and Liang, 2021) insert learnable vectors into the input or attention layers, and adapter tuning (Chen et al., 2022) adds lightweight modules inside transformer layers. While effective in many domains, these approaches often struggle with the multi-scale structures, rich spectral content, and complex semantics of RS data. To bridge this gap, RS-specific PEFT designs have been proposed: AiRs (Hu et al., 2024b) models spatial context and semantic responses, TEA (Hu et al., 2024a) employs a lightweight side adapter with macro-scale semantic guidance, and Earth-Adapter (Hu et al., 2025) applies frequency decoupling for multi-scale modeling but relies on a unified linear encoder, limiting scale-specific feature modeling.
@@ 2.2. Parameter-Efficient Adaptation for Multi Scale Remote Sensing Understanding
To address challenges in remote sensing imagery, we propose the FreqWeaver Adapter, a novel adapter design that integrates frequency-domain decomposition with convolutional processing. It separates high- and low-frequency components through dedicated convolutional branches and incorporates spatial information. This design enables efficient multi-scale adaptation, reduces aliasing artifacts, avoids large variations in object scale, while maintaining a very low additional parameter cost.
2.3. Semi-Supervised Learning for Robust Remote Sensing Representations
Semi-supervised learning (SSL) leverages both labeled and unlabeled data to improve model generalization and robustness under limited annotations. General SSL strategies include pseudo-labeling (Lee et al., 2013; Xie et al., 2020), consistency regularization (Berthelot et al., 2019; Sohn and Berthelot, 2020), and teacher student frameworks (Tarvainen and Valpola, 2017). While these approaches have demonstrated strong performance in natural image tasks, their direct application to remote sensing (RS) is often limited by challenges such as multi-scale spatial structures and high inter-class spectral similarity. To overcome these limitations, several methods have extended SSL to better fit RS characteristics: SSSM (Feng et al., 2025) introduced multi-level feature fusion with pseudo-label sampling for scene classification, alleviating the noise caused by low-confidence labels; UniMatch (Yang et al., 2023) enhanced semantic segmentation with task-driven perturbation branches to better handle complex RS structures; MUCA (Wang et al., 2025) incorporated multi-scale uncertainty consistency and cross teacher–student attention to align intermediate features across scales, improving the discrimination of spectrally similar land-cover classes; and SOOD (Hua et al., 2023) adapted semi-supervised learning to oriented object detection by introducing orientation-consistency and global-consistency losses, significantly improving detection performance on aerial benchmarks. Overall, these methods highlight the potential of SSL to reduce annotation costs and enhance performance in complex RS scenarios, but their applications to UHSR data remain insufficiently explored.
To reduce the noise caused by unstable predictions in UHSR imagery, we propose the Uncertainty-Aware Teacher–Student Framework. This framework uses entropy-based uncertainty estimation to select high-confidence pseudo-labels and focuses supervision on reliable regions. In this way, it limits the effect of noisy predictions, makes better use of large amounts of unlabeled data, and learns more stable feature representations while lowering dependence on extensive annotations.
Figure 1. Overview of the proposed framework. Given a remote sensing image, two augmented views are generated: a strongly augmented version is fed into the student network, and a weakly augmented version is processed by the teacher network. To capture the characteristics of UHSR data, a parameter-efficient FreqWeaver Adapter is inserted into specific Transformer layers. The teacher network is updated with an exponential moving average (EMA) strategy to maintain stable parameter evolution. For unlabeled data, the teacher network provides pseudo-labels to enforce consistency with student predictions, enabling more robust feature learning. For labeled data, ground-truth annotations directly supervise the student network to ensure accurate semantic understanding. Overall, the framework integrates adapter-based adaptation with semi-supervised learning to deliver reliable segmentation of UHSR imagery while reducing reliance on labeled data.